# Time Matters: Scaling Laws for Any Budget

### v1

@@ Abstract
A primary cost driver for training large models is wall-clock training time. We show that popular time estimates based on FLOPs are poor estimates, and construct a more accurate proxy based on memory copies. We show that with some simple accounting, we can estimate the training speed of a transformer model from its hyperparameters. Combined with a scaling law curve like Chinchilla, this lets us estimate the final loss of the model. We fit our estimate to real data with a linear regression, and apply the result to rewrite Chinchilla in terms of a model’s estimated training time as opposed to the amount of training data. This gives an expression for the loss in terms of the model’s hyperparameters alone. We show that this expression is accurate across a wide range of model hyperparameter values, enabling us to analytically make architectural decisions and train models more efficiently.
@@ 1 Introduction
2 The parameter equivalence principle
The core intuition motivating this work is the observation that large models are not particularly sensitive to their hyperparameters, provided we hold the total parameter count constant. This idea was discussed in [1], but due to its importance we capture it in a form of an equivalence principle.
Principle 1 (The Parameter Equivalence Principle).
Above a certain scale, the final loss of a transformer is primarily a function of how many parameters there are, not where they are in the model.
@@ 3 Estimating linear scaling law Coefficients
Hoffmann et al. [2] and Besiroglu et al. [4] derive different coefficient values with the most extreme difference being their linear data coefficient $B$. Rather than enter into this debate, we simply take the exponents from [2], and fit our own linear coefficients $A$, $B$ and $E$ using linear regression on the model loss. This was done by iterating over 1,535 different decoder-only transformer models’ hyperparmeters configurations trained from scratch for three hours each on the C4 Dodge et al. [3] dataset. Models sizes vary from $319K$ to $310M$ parameters. Note that we constrained our experiments to models that can be trained on a single TPU to avoid confounders from inter-chip communication. We experimented with model hyperparameters of embed sizes ranging from $2^{5}$ to $2^{10}$, number of layers ranging from $3$ to $8$, MLP width ranging from $2^{8}$ to $2^{14}$, number of attention heads ranging from $2^{1}$ to $2^{7}$, and a fixed vocabulary size of 8,000. We trained on a mesh of 4 hosts x 8 chips/host of TPU V5 chips with no model sharding.
@@ 5 Estimating the throughput
We conduct a large scale ($N=3,556$) sweep over model hyperparameters trained for 5 minutes, just long enough to accurately determine the number of tokens per second they process. We applied linear regression over the data to determine $c_{1}$, $c_{2}$, and $c_{3}$.
@@ 6 Putting it all together
We now have an equation that estimates the number of tokens that the model will consume from its hyperparameters. We also have an exact expression for the number of parameters in such a model, PARAMS. Our tuned Chinchilla equation relates these two quantities to estimate the final loss (3). In figure 4, we show the results of this estimation, applied to the data from Section 3.
@@
Notice that the graph is largely indistinguishable from Figure 1, including the quality of fit $r^{2}=0.9$. While there is some error in the Chinchilla equation’s predictions, there is essentially no additional error from using our estimates in place of the empirical values.
@@ 7 Better loss with faster models
We can see that increasing the embed size at the expense of the other hyperparameters is favorable throughout the plotted region. hidden_size is particularly disincentivized. This suggests that we should take our MLPs to be narrow, and our models to be somewhat shallow, in exchange for much larger embed size.
@@ 8 Conclusion
We demonstrated this effect across a wide variety of model sizes, and showed that we can accurately predict the model’s loss from its hyperparameters, simply by estimating how many memory copies will take place. Crucially, this is a stronger predictor than approaches based on FLOPs. What’s more, it allows us to make specific predictions about which hyperparameters to use during model design.

### v2

@@ Abstract
A primary cost driver for training large models is wall-clock training time. We show that popular time estimates based on FLOPs are poor estimates, and construct a more accurate proxy based on memory copies. This allows us to accurately estimate the training speed of a transformer model from its hyperparameters. Combined with a scaling law curve like Chinchilla, this allows us to accurately predict the final loss of a model from a simple equation. We show that this expression is accurate across a wide range of model hyperparameter values, enabling us to analytically make architectural decisions and train models more efficiently. Crucially, this analysis predicts that in contrast to existing literature, models should be wider rather than deeper, as the benefits of speed outweigh the benefits of depth.
@@ 1 Introduction
2 The parameter equivalence heuristic
The core intuition motivating this work is the observation that large models are not particularly sensitive to their hyperparameters, provided we hold the total parameter count constant. This idea was discussed in [1], but due to its importance we capture it in a form of an equivalence heuristic.
Principle 1 (The Parameter Equivalence Heuristic).
Above a certain scale, the final loss of a transformer at the end of training is primarily a function of how many parameters there are, not where they are in the model.
@@ 3 Estimating linear scaling law Coefficients
Hoffmann et al. [2] and Besiroglu et al. [4] derive different coefficient values with the most extreme difference being their linear data coefficient $B$. Rather than enter into this debate, we simply take the exponents from [2], and fit our own linear coefficients $A$, $B$ and $E$ using linear regression on the model loss. This was done by iterating over 767 different decoder-only transformer models’ hyperparmeters configurations trained from scratch for three hours each on the C4 Dodge et al. [3] dataset and then evaluating our predictions over a holdout set of 767 different model configurations trained in the same manner over the same dataset. Models sizes vary from $319K$ to $310M$ parameters. Note that we constrained our experiments to models that can be trained on a single TPU to avoid confounders from inter-chip communication. We experimented with model hyperparameters of embed sizes ranging from $2^{5}$ to $2^{10}$, number of layers ranging from $3$ to $8$, MLP width ranging from $2^{8}$ to $2^{14}$, number of attention heads ranging from $2^{1}$ to $2^{7}$, and a fixed vocabulary size of 8,000. We trained on a mesh of 4 hosts x 8 chips/host of TPU V5 chips with no model sharding.
@@ 5 Estimating the throughput
We conduct a larger scale ($N=1,778$) sweep over model hyperparameters trained for 5 minutes, just long enough to accurately determine the number of tokens per second they process. We applied linear regression over a holdout set of the same size($N=1,778$) to determine that $c_{1}=3.74e-19$, $c_{2}=2.4e-15$, and $c_{3}=1.46e-07$.
@@ 6 Putting it all together
We now have an equation that estimates the number of tokens that the model will consume from its hyperparameters. We also have an exact expression for the number of parameters in such a model, PARAMS. Our tuned Chinchilla equation relates these two quantities to estimate the final loss (3). In figure 4, we show the results of this estimation, applied to the holdout data from Section 3.
@@
Notice that the graph is largely indistinguishable from Figure 1, including the quality of fit $r^{2}=0.92$. While there is some error in the Chinchilla equation’s predictions, there is essentially no additional error from using our estimates in place of the empirical values.
@@ 7 Better loss with faster models
We can see that increasing the hidden size at the expense of the other hyperparameters is favorable throughout the plotted region. num_layers is particularly disincentivized. This suggests that in contrast with common practice we should take our MLPs to be wide, and our models to be somewhat shallow, in exchange for smaller embed size.
8 Other Architectures
There are many architectures other than transformers for which this kind of technique could be of interest. We focused on transformers as they are the dominant architecture used in practice today. With that said, our core observations are simply that
  • If there exists a scaling law that can predict loss in terms of the number of parameters and the amount of data, the lowest loss will be achieved when we arrange the parameters to make the model as fast as possible, allowing us to train on more data.
  • Estimating speed correctly requires considering not only the FLOPs but also the MEMCPYs.
  • FLOPs, MEMCPYs and the total number of parameters can all be counted using simple accounting.
These observations are not specific to transformers, even if the particulars of the accounting differs. For this reason, we expect analogous results for other architectures.
@@ 9 Conclusion
We demonstrated this effect across a wide variety of model sizes, and showed that we can accurately predict the model’s loss from its hyperparameters, simply by estimating how many memory copies will take place. Crucially, this is a stronger predictor than approaches based on FLOPs. In particular, it predicts that in contrast with common practice we should be making shorter, wider models, because the speed benefits dominate in practice.